Hi! I have a question that asks me to find the equation of a line with gradients and intercepts on the y axis. I am given points such as -2/7, 6/7. With those two points I understand the answer is 7y+6-2x. But I can't get my head round the answer to this one following (and my exams are sooo soon!): Can you help me to understand -4/5, 2 and why this gives the answer 5y+4x=10?
Thanks ever so much!
~Chris

Answer by stanbon(75887)

I have a question that asks me to find the equation of a line with gradients and intercepts on the y axis. I am given points such as -2/7, 6/7.
slope = -2/7 ; y-intercept = 6/7
y = (-2/7)x + (6/7)
7y = -2x + 6
7y =
With that information I understand the answer is 7y = 6-2x.
------------------------
But I can't get my head round the answer to this one following (and my exams are sooo soon!): Can you help me to understand -4/5, 2 and why this gives the answer 5y+4x=10?
------------------
slope = -4/5
y-intercept = 2
--------------------------
y = (-4/5)x + 2
5y = -4x + 10
5y + 4x = 10
===================
Cheers,
Stan H.